# Linearly-scalable and entropy-optimal learning of nonstationary and nonlinear manifolds

### v2

@@ 4 Visualising EOMC internal coordinates $\gamma$ with t-SNE
Figure 8: Nonlinear t-SNE for visualisation of the EOMC internal coordinates $\gamma$ from (2.2). Panel A: visualising $\gamma$ from the Example 1 in Sec. 3.1.1 (left), and $\gamma$ from the Example 2 in Sec. 3.1.2. Panel B: t-SNE visualisation of raw data from the L96 model described in Sec. 3.2 (left column), and for EOMC $\gamma$ (right column).
@@ 5 Discussion
In contrast, as was analysed above in Sec. 3.2, metastability of $\gamma$ can not be an artefact of the EOMC data analysis: as can be seen from (2.2), the EOMC loss function does not contain any terms that would enforce the persistence, metastability or Markov-property on $\gamma$. As a matter of fact, the value of $\mathcal{L}^{\textrm{EOMC}}$ in (2.2) is invariant with respect to any permutation of the columns of data matrix $X$. This means that the observed metastability of $\gamma$ for Lorenz-96 in chaotic and very chaotic regimes can only be an imprint of the underlying L96 dynamics - that appears to be best described by a metastable process switching between low-dimensional (with $d=3$) locally-linear manifolds. Analysis from Sec. 3.2 revealed that the main effect of increasing forcing was in gradually-increasing the total number $K$ of these local low-dimensional manifolds - and in slowly decreasing mean relaxation and mean exit times, that were measured from the transfer operators inferred from EOMC variables $\gamma$ (see Fig. 6). We shown that average Markovian mean exit times $\mathbb{E}_{k}\left[\tau_{k}^{ex}(P)\right]$ closely match the doubling times from positive Lyapunov exponents in the literature (see Fig. 6B) - indicating the predictability horizons quantified with Lyapunov exponents as limits for the average times it takes for the system to leave the current low-dimensional manifold and to go somewhere else. And, it was shown that the transfer operator description revealed by the EOMC analysis in Figs. 5 and 6 allows going beyond this limit: besides allowing to measure the average time $\tau_{k}^{ex}(P)$ the dynamics spends in the local manifold $k$, transfer operator provides the transition probabilities $P_{j,k}$ to all of the other local manifold states $j\neq k$, where the dynamics can go to after leaving $k$. For Markov processes, relaxation times define the predictability horizon of the system, quantifying the time that the system requires to forget its initial condition [34]. As revealed by the relaxation times curve $\tau^{rel}(P)$ (see the blue dotted curve in the Fig. 6B), this transfer operator description roughly doubles the prediction horizon for the system, as compared to the state of the art Lyapunov exponent description.

### v3

@@ 4 Visualising EOMC internal coordinates $\gamma$ with t-SNE
Figure 8: Nonlinear t-SNE for visualisation of the EOMC internal coordinates $\gamma$ from (17-20). Panel A: visualising $\gamma$ from the Example 1 in Sec. 3.1.1 (left), and $\gamma$ from the Example 2 in Sec. 3.1.2. Panel B: t-SNE visualisation of raw data from the L96 model described in Sec. 3.2 (left column), and for EOMC $\gamma$ (right column).
@@ 5 Discussion
In contrast, as was analysed above in Sec. 3.2, metastability of $\gamma$ can not be an artefact of the EOMC data analysis: as can be seen from (17-20), the EOMC loss function does not contain any terms that would enforce the persistence, metastability or Markov-property on $\gamma$. As a matter of fact, the value of $\mathcal{L}^{\textrm{EOMC}}$ in (17-20) is invariant with respect to any permutation of the columns of data matrix $X$. This means that the observed metastability of $\gamma$ for Lorenz-96 in chaotic and very chaotic regimes can only be an imprint of the underlying L96 dynamics - that appears to be best described by a metastable process switching between low-dimensional (with $d=3$) locally-linear manifolds. Analysis from Sec. 3.2 revealed that the main effect of increasing forcing was in gradually-increasing the total number $K$ of these local low-dimensional manifolds - and in slowly decreasing mean relaxation and mean exit times, that were measured from the transfer operators inferred from EOMC variables $\gamma$ (see Fig. 6). We shown that average Markovian mean exit times $\mathbb{E}_{k}\left[\tau_{k}^{ex}(P)\right]$ closely match the doubling times from positive Lyapunov exponents in the literature (see Fig. 6B) - indicating the predictability horizons quantified with Lyapunov exponents as limits for the average times it takes for the system to leave the current low-dimensional manifold and to go somewhere else. And, it was shown that the transfer operator description revealed by the EOMC analysis in Figs. 5 and 6 allows going beyond this limit: besides allowing to measure the average time $\tau_{k}^{ex}(P)$ the dynamics spends in the local manifold $k$, transfer operator provides the transition probabilities $P_{j,k}$ to all of the other local manifold states $j\neq k$, where the dynamics can go to after leaving $k$. For Markov processes, relaxation times define the predictability horizon of the system, quantifying the time that the system requires to forget its initial condition [34]. As revealed by the relaxation times curve $\tau^{rel}(P)$ (see the blue dotted curve in the Fig. 6B), this transfer operator description roughly doubles the prediction horizon for the system, as compared to the state of the art Lyapunov exponent description.